# Additive Sink Subtraction

### v1

@@ Abstract
Subtraction games are a classical topic in Combinatorial Game Theory. A result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length. Flammenkamp (1996) conjectures a striking classification for three-move subtraction games: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths. Despite early attention in Winning Ways (1982), the general additive case remains open. We introduce and analyze a dual winning convention, which we call sink subtraction. Unlike the standard wall convention, where moves to negative positions are forbidden, the sink convention declares a player the winner upon moving to a non-positive position. We show that additive sink subtraction admits a complete solution: the nim-sequence is purely periodic with an explicit linear or quadratic period formula, and we conjecture a duality between additive sink subtraction and classical wall subtraction. Keywords: Additive Subtraction Game, Nimber, Periodicity, Sink Convention.
@@ 1 Introduction
We study classical combinatorial subtraction games [1], with a twist on the terminal condition. Let $S\subset\mathbb{N}=\{1,2,\ldots\}$. For any $x\in\mathbb{N}$ and all $s\in S$, $x-s\in\mathbb{Z}$ is a move option. The sink is an infinite sequence of terminal P-positions for all non-positive integers. No move is possible from a position in the sink. We study two-player impartial games in the normal-play convention, so “a player who cannot move loses”, and this is the same as “a player who plays into the sink wins”. We denote this ruleset family by sink subtraction. Observe that, if $S=\mathbb{N}$, we play the game of nim and sink subtraction coincides with wall subtraction (our terminology), where all negative heap sizes are forbidden. In Figure 1, we illustrate with the ruleset $S=\{1,5\}$ how sink vs. wall play can differ.
Figure 1: sink and wall subtraction, respectively. Suppose the ruleset is $S=\{1,5\}$. To the Left, the boy can choose to move into the sink or he can play to the tip of the jetty, and let daddy win. To the Right, the boy has only one move, because the larger move is prohibited by the wall. Thus the wall-game is determined already by parity.
@@
Indeed, if $S$ is finite, the two ruleset families sink and wall usually satisfy distinct nimber patterns. A game’s nimber (a.k.a. nim-value) is defined via the standard minimum exclusive rule; for any finite $X\subset\mathbb{N}_{0}$, let $\mathrm{mex}X=\min\{\mathbb{N}_{0}\setminus X\}$, where $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. Then the nim value algorithm of $x$ is $v(x)=\mathrm{mex}\{v(x^{\prime})\}$, where $x^{\prime}$ ranges over the options of $x$. Note that this gives value zero for all positions in the sink.
@@
with a pre-period of size three but the same period length 7. The periodic patterns are the same but occur in a different order: 0011021 and 2100110 respectively. To the best of our knowledge, sink subtraction has not yet been studied in the literature. It is the purpose of this note to enter this unpaved territory. We use the convention of $x=0$ belonging to the sink, so we start nim-sequences by the (non-zero) value of position one.
Of course, the folklore periodicity theorem holds for rulesets where the terminal condition is a sink instead of a ‘wall’, with exactly the same proof.
@@
The purpose of this note is to establish the period lengths of the ruleset family additive sink subtraction $S$ with subtraction set
@@
where $m,\delta\in\mathbb{N}$. There is a conjecture from Flammenkamp’s PhD thesis [2] concerning the periodicity of Additive Wall Subtraction, and this is to some degree paraphrasing the treatment in Winning Ways [1], where the linear-period-length part of the conjecture has been proved. As far as the authors know, the non-linear (quadratic) part has not yet been established. [^1] Non-additive three-move games are conjectured to have period lengths of the form “the sum of two moves”, where “which two moves?” appear to satisfy some fractal patterns [2]. Let us state our main result.
@@
Let $m,\delta\in\mathbb{N}$. Consider additive sink subtraction with subtraction set $S(m,\delta)=\{m,m+\delta,2m+\delta\}$, and let $d:=\delta\ \pmod{2m}$, with $0\leq d<2m$. Then $S$ is purely periodic, and its period length is
@@ 2 A proof of the main result
Consider item (ii). For convenience we restrict $m<\delta<2m$, and let $\delta=m+k$. (The general case turns out to be an immediate consequence.) As an example, let us begin by describing the periodic word for any rulesets of the form “the lower case”, i.e. whenever $k=1$. The word is described by consecutive “$B$-blocks” of the form
@@
for $0\leq i\leq m-1$, and at last, we concatenate the factor $0^{m}$. By increasing the ‘block speed’, the factor $1\leq k<m$, another type of blocks appears, the “$C$-blocks”, and the extreme “upper case” will favor only the $C$-blocks. The challenge is to understand the intermediate cases, with mixing between the block types; in Figure 2, we illustrate all periods for the case of $m=19$, $0<k<19$.
Figure 2: Nimber periods for $m=19$ and $0<k<19$, starting from $x=1$. The colors correspond to the nimbers; purple (0), blue (1), green (2) and yellow (3). Zoom in for details.
@@
Let us prove that the construction produces the correct period lengths. For example, if $m=5$ and $k=1$, then (by ignoring technical details) the block structure is $BBBBB0^{m}$, which satisfies the claimed period length $5(4\cdot 5+2c-2a)+5=115$. If $m=5$ and $k=4$, then the block structure is $BCCCB0^{m}$, which satisfies the claimed period length $2(4\cdot 5+8)+3(6\cdot 5+4-b+a)+5=56+99+5=160$. The cases for $m=5$ are illustrated in the upper picture of Figure 3. If $m=6$ and $k=2$, then the block structure is $BBB0^{m}BBB0^{m}$, with claimed period length $3(4\cdot 6+2c-2a)+6=90$. If $m=6$ and $k=4$, then the block structure is $BCB0^{m}BCB0^{m}$, with claimed period length $2(4\cdot 6+2c-2a)+6\cdot 6+4+b-a+6=108$. The cases for $m=6$ are illustrated in the lower picture of Figure 3.
Recall that we here assume $\delta=m+k$, $1\leq k<m$.
@@
The blocks use two independent indices; the second one will be used throughout the tables to come.
@@
    Within a fixed block, we label its consecutive factors by a local index. For a $B$-block, the factors are denoted $\texttt{B1},\ldots,\texttt{B5}$, where Bj refers to the (j+3)(j+3)rd factor of that block, for $j\in\{1,\ldots,5\}$. Similarly, Aj denotes the $j$th factor of the prefix $A(i)$ associated with the same block index $i$.
Unless explicitly stated otherwise, the block index $i$ is fixed when referring to factor indices. For example, the notation B3 always refers to the sixth factor within a single $B(i)$ block, independent of the block index $i$. Whenever we refer to an object other than a block or a factor, we explicitly use the term word.
Recall that the block length of $B$ is $4m-2\alpha+2\gamma=4m+2k>3m+k$ (by $\gamma>\alpha$), while the block length of $C$ is $6m+k-\alpha+\beta=5m+2k>3m+k$ (by $\beta<\alpha$). One can check that, under block concatenation, the distance between two similar factors is always greater than $3m+k$; two A factors are similar, if they carry the same j index, and analogously for B and C.
We will treat each factor as a single entity, and use terminology such as factor-option and factor value. The advantage is that it drastically reduces the number of cases, but an apparent disadvantage is that sometimes individual positions within a factor can have options in distinct factors, for a given move. This phenomenon proves unavoidable though, and is part and parcel of the quadratic period length, as we will see. By intuition, when factor relations settle, then we approach the full period.
We use elementary algebra to justify mex-consistency. Each factor $F$, of $P$, is of the form $F=x^{y}$ with $x\in\{0,1,2,3\}$ and where $y\geq 0$ is maximum such that adjacent letters differ from $x$; the length of $F$ is $|F|=y$. The factor value of a factor $F$ is $v(F)=x$ if $F=x^{y}$. We often want to identify a factor with its corresponding set of positions. To this purpose it is convenient to temporarily identify position one with the first letter of some relevant factor, and we write F to denote the factor set, the set of such induced positions corresponding to a factor $F$. The value of a factor set is its factor value. A factor-option of $F$ is a factor $F^{\prime}$ such that there is an $x\in\texttt{F'}$ such that, for some $y\in\texttt{F}$, $x-y\in\{m,2m+k,3m+k\}$. Any such terminology is local and used as a computational device.
@@
Throughout the tables, the symbols $s_{1}$, $s_{2}$, and $s_{3}$ denote the moves $m$, $2m+k$, and $3m+k$, respectively. In all reachability tables, a dash “–” indicates that no reachability condition is required from the corresponding factor to that value. In tables for concatenations (e.g. BB), we use uppercase factor names for the current block $B(i)$ and lowercase for the preceding block B(i1)B(i-1). For example, B4 is the set of positions corresponding to the factor $0^{m-\alpha}$ in $B(i)$, while b4 is the set of positions corresponding to the factor $0^{m-\alpha^{\prime}}$ in B(i1)B(i-1); Throughout, we write $\alpha^{\prime}=\alpha(i-1)$, $\beta^{\prime}=\beta(i-1)$ and $\gamma^{\prime}=\gamma(i-1)$.
In the tables, rather than fixing a specific heap of size one, we adopt the following convention. For each row (i.e. each game value under consideration), we choose as reference point the smallest position in the leftmost factor of that value; for the sink we use position 0 as reference. Each factor is then interpreted as a set of consecutive positions, all carrying the same game value, determined relative to this reference point. In this way, the symbolic word representations, together with the specified factor lengths, are mapped unambiguously to actual game positions, allowing us to analyze move options between factors.
In the reachability tables, distances are measured from the beginning of the relevant factor to the end of the preceding block. For the factor b4, the distance to the end of the $B$-block is $d(\texttt{b4})=m-\gamma^{\prime}+\alpha^{\prime}=m-k$ (by Lemma 5), and its length is $|\texttt{b4}|=\gamma(i-1)=\gamma^{\prime}$. For the relevant $c$-factors, the distances to the end of the $C$-block are $d(\texttt{c4})=2m+k\text{ and }d(\texttt{c8})=m+\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}\text{ and }|\texttt{c8}|=m.$
Throughout, “factor” refers to a symbolic word of the form $x^{y}$, while “factor set” refers to the induced set of game positions when a reference point is fixed. We only pass to factor sets when verifying reachability or anti-collision. To read the tables: fix a column factor set. Each row lists all potential factor-options of a relevant game value. Reachability tables ensure the existence of smaller values as move options, while anti-collision tables exclude equal values in the set of move options; here a triple of inequalities corresponds to the moves $s_{1},s_{2},s_{3}$, respectively.
Example A3 in bZB. This is the case where $\beta(i-1)=\alpha(i)=0$, and hence $\zeta(i-1)=0^{m}$. Moreover, it follows that $\alpha(i-1)=m-k$. The factor set A3 has size $m$ in such $B(i)$, and, by (14), we study the word
@@ 3 Final remarks
On a more philosophical matter, with respect to the ease of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games. However this type of result has been on the list of conjectures for a long time, and our conjectured reciprocity with the wall convention requires further research. For example, Richard Guy writes in the first problem, in the first edition of Unsolved problems in combinatorial games [4] “Subtraction games are known to be periodic. Investigate the relationship between the subtraction set and the length and structure of the period.”; see also [6]. The second part of the conjecture for three move games [2, 7] concerns the non-additive three-move rulesets. Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [2], and has not yet been fully understood. See also the recent survey on subtraction games [5].
Acknowledgements. Thanks to ChatGPT 5.2 for the two images in Figure 1.

### v2

@@ Abstract
Subtraction games are a classical topic in Combinatorial Game Theory. An early result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length. Even three-move games are not yet fully understood. Nevertheless, Flammenkamp (1997) conjectures a striking classification: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths. Despite early attention in Winning Ways (1982), the general additive case remains open. We analyze a dual winning convention, called sink subtraction. Unlike the standard wall convention, where moves to negative positions are forbidden, the sink convention declares a player the winner upon moving to a non-positive position. We show that additive sink subtraction admits a complete solution: its nim-sequence is purely periodic with an explicit linear or quadratic period formula. We further conjecture a nimber duality between additive sink subtraction and classical wall subtraction.
Keywords: additive subtraction game, nimber, periodicity, sink convention.
@@ 1 Introduction
We study classical combinatorial subtraction games [2], with a twist on the terminal condition. Let $S\subset\mathbb{N}=\{1,2,\ldots\}$. For any $x\in\mathbb{N}$ and all $s\in S$, $x-s\in\mathbb{Z}$ is a move option. The sink is an infinite sequence of terminal $\mathcal{P}$-positions ($\mathcal{P}$revious player winning positions) for all non-positive integers. No move is possible from a position in the sink. We study two-player impartial games in the normal-play convention, so “a player who cannot move loses”, and this is the same as “a player who plays into the sink wins”. We denote this ruleset family by sink subtraction. Observe that, if $S=\mathbb{N}$, we play the game of nim and sink subtraction coincides with wall subtraction (our terminology), where all negative heap sizes are forbidden. In Figure 1, we illustrate with the ruleset $S=\{1,5\}$ how sink vs. wall play can differ.
Figure 1: sink and wall subtraction, respectively. Suppose the ruleset is $S=\{1,5\}$. To the Left, the boy can choose to move into the sink or he can play to the tip of the jetty, and let daddy win. To the Right, the boy has only one move, because the larger move is prohibited by the wall. Here, the wall game is determined already by parity.
@@
Indeed, if $S$ is finite, the two ruleset families sink and wall usually satisfy distinct nimber patterns. A game’s nimber (a.k.a. nim-value) is defined via the standard minimum exclusive rule; for any finite $X\subset\mathbb{N}_{0}$, let $\mathrm{mex}X=\min\{\mathbb{N}_{0}\setminus X\}$, where $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. Then the nim-value algorithm of $x$ is $v(x)=\mathrm{mex}\{v(x^{\prime})\}$, where $x^{\prime}$ ranges over the options of $x$. Note that this gives value zero for all positions in the sink.
@@
with a pre-period of size three but the same period length 7. The periodic patterns are the same but occur in a different order: 0011021 and 2100110 respectively.
To the best of our knowledge, sink subtraction has only been studied in the literature in terms of outcomes, and only in one paper [1]. They remark that one can relate the outcomes of the two conventions via the formula
$o_{\rm wall}(x)=o_{\rm sink}(x+\max(S)+1).$ (1)
[…]
In this paper we continue the pursuit of [1], by studying the more general nim-values of sink subtraction. We use the convention of position $x=0$ belonging to the sink, so we start nim-sequences by the value (one) of position one. Of course, the folklore periodicity theorem holds for rulesets where the terminal condition is a sink instead of a ‘wall’, with exactly the same proof.
@@
The purpose of this paper is to establish the period lengths of the ruleset family additive sink subtraction $S$ with subtraction set
@@
where $m,\delta\in\mathbb{N}$. There is a conjecture from Flammenkamp’s PhD thesis [3] concerning the periodicity of additive wall subtraction, and this partially paraphrases the treatment in Winning Ways [2], where the linear-period-length part of the conjecture has been proved. As far as the authors know, the non-linear (quadratic) part has not yet been established. [^1] Non-additive three-move games are conjectured to have period lengths of the form “the sum of two moves”, where “which two moves?” appear to satisfy some fractal patterns [3]. Let us state our main result.
@@
Let $m,\delta\in\mathbb{N}$. Consider additive sink subtraction with subtraction set $S(m,\delta)=\{m,m+\delta,2m+\delta\}$, and let $d:=\delta\ \pmod{2m}$, with $0\leq d<2m$. Then $S$ has purely periodic nim-values, with period length
@@
The immediate consequence for the outcomes of additive wall subtraction is as follows.
Corollary 4.
Let $m,\delta\in\mathbb{N}$, and consider additive wall subtraction with subtraction set $S(m,\delta)=\{m,m+\delta,2m+\delta\}$. Then the period length of the outcomes is the same as in Theorem 3, $p(m,\delta)$.
Proof.
This is obvious by (1). ∎
We conjecture that there is an explicit duality formula for relating the nimbers of additive sink subtraction with those of additive wall subtraction, of the form, for all $x$,
$v_{\rm wall}\bigl([x]_{p}\bigr)=\sigma\bigl(v_{\rm sink}\bigl([s_{2}-x]_{p}\bigr)\bigr),$ (3)
where $\sigma:0\rightarrow 2,1\rightarrow 1,2\rightarrow 0,3\rightarrow 3$. Here $x]_{p}$ denotes the heap size modulo the (outcome) period length $p$.
Example 5.
The wall and sink nim-sequences of $S=\{2,5,7\}$ are displayed in the below table; they have period length $22$. The upper (lower) row represents the heap sizes for wall (sink), and the middle row entries illustrate the respective nimbers (wall upper and sink lower). Note that the values satisfy Equation 3.
$x_{\mathrm{wall}}$ 0 11 22 33 44 55 66 77 88 99 $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ 1919 2020 2121
$v(x)$ $0\atop 2$ $0\atop 2$ 11 11 $0\atop 2$ $2\atop 0$ 11 33 $2\atop 0$ $2\atop 0$ $0\atop 2$ 33 11 $0\atop 2$ $0\atop 2$ 11 11 $2\atop 0$ $2\atop 0$ 33 33 $2\atop 0$
$x_{\mathrm{sink}}$ $17$ $16$ $15$ $14$ $13$ $12$ $11$ $10$ 99 88 77 66 55 44 33 22 11 $22$ 2121 2020 1919 $18$
@@ 2 A proof of the main result
Consider item (ii). For convenience we restrict $m<\delta<2m$, and let $\delta=m+k$. (The general case turns out to be an immediate consequence.) Similar to case (i), let the alphabet be $\{0,1,2,3\}$. As an example, let us begin by describing the periodic word for any rulesets of the form “the lower case”, i.e. whenever $k=1$. The word is described by consecutive “$B$-blocks” of the form
@@
for $0\leq i\leq m-1$, and at last, we concatenate the factor $0^{m}$. [^2] The period length for this case is mentioned in [1], but without giving the explicit block construction or a proof. By increasing the ‘block speed’, the factor $1\leq k<m$, another type of blocks appears, the “$C$-blocks”, and the extreme “upper case” will favor only the $C$-blocks. The challenge is to understand the intermediate cases, with mixing between the block types; in Figure 2, we illustrate all periods for the case of $m=19$, $0<k<19$.
Figure 2: Nimber periods for $m=19$ and $0<k<19$, starting from $x=1$. The colors correspond to the nimbers; purple (0), blue (1), green (2) and yellow (3). The rows have different scaling as to fit the full periods in each case. Zoom in for details.
@@
Let us prove that the construction produces the correct period lengths. For example, if $m=5$ and $k=1$, then (by ignoring technical details) the block structure is $BBBBB0^{m}$, which satisfies the claimed period length $5(4\cdot 5+2c-2a)+5=115$. If $m=5$ and $k=4$, then the block structure is $BCCCB0^{m}$, which satisfies the claimed period length $2(4\cdot 5+8)+3(6\cdot 5+4-b+a)+5=56+99+5=160$. The cases for $m=5$ are illustrated in the upper picture of Figure 3. If $m=6$ and $k=2$, then the block structure is $BBB0^{m}BBB0^{m}$, with claimed period length $3(4\cdot 6+2c-2a)+6=90$. If $m=6$ and $k=4$, then the block structure is $BCB0^{m}BCB0^{m}$, with claimed period length $2(4\cdot 6+2c-2a)+6\cdot 6+4+b-a+6=108$. The cases for $m=6$ are illustrated in the middle picture of Figure 3. Recall that we here assume $\delta=m+k$, $1\leq k<m$.
@@
We use elementary algebra to justify mex-consistency. Each maximum factor $F$, of $P$, is of the form $F=a^{b}$ with $a\in\{0,1,2,3\}$ and where $0\leq b\leq m$ is maximum such that adjacent letters differ from $a$; the length of $F$ is $|F|=b$. We often omit the word ‘maximum’ and call a maximum factor simply by a factor. The factor-value of a factor $F$ is $v(F)=a$ if $F=a^{b}$. We will treat each factor as a single entity, and use terminology such as factor-option and factor-value. The advantage is that it drastically reduces the number of cases, but an apparent disadvantage is that sometimes individual positions within a factor can have options in distinct factors, for a given move. This phenomenon proves unavoidable though, and is part and parcel of the quadratic period length, as we will see. By intuition, when factor relations settle, then we approach the full period.
The blocks will use two independent indices, the latter of which will enumerate its factors.
@@
    Within a fixed block, we label its consecutive factors by a local index. For a given $B(i)$-block, the factors are denoted $A_{1},A_{2},A_{3},B_{1},\ldots,B_{5}$, where $A_{j}$ denotes the $j$th factor of the prefix $A(i)$ and $B_{j}$ refers to the (j+3)(j+3)rd factor of that block.
Unless explicitly stated otherwise, the block index $i$ is fixed when referring to factor indices. For example, the notation $B_{3}$ always refers to the sixth factor within a single $B(i)$ block, where the block index $i$ is given by the surrounding context. Whenever we refer to an object other than a block or a factor, we explicitly use the term word.
Recall that the block length of $B$ is $4m-2\alpha+2\gamma=4m+2k>3m+k$ (by $\gamma>\alpha$), while the block length of $C$ is $6m+k-\alpha+\beta=5m+2k>3m+k$ (by $\beta<\alpha$). One can check that, under block concatenation, the distance between two similar factors is always greater than the largest move, $3m+k$; two $A$-factors are similar, if they carry the same $j$-index, and analogously for $B$- and $C$-factors.
@@
We often study the corresponding set of positions of a given factor. To this purpose it is convenient to temporarily identify position one with the first letter of some relevant factor, and we use a special font F to denote the factor set, the set of such induced positions corresponding to a factor $F$. Any such notation is local and used as a computational device. Each factor is then interpreted as a set of consecutive positions, all carrying the same game value, determined relative to this reference point. In this way, the symbolic word representations, together with the specified factor lengths, are mapped unambiguously to actual game positions, allowing us to analyze move options between factors.
Throughout the tables, the symbols $s_{1}$, $s_{2}$, and $s_{3}$ denote the moves $m$, $2m+k$, and $3m+k$, respectively. In all reachability tables, a dash “–” indicates that no reachability condition is required from the corresponding factor to that value. In tables for concatenations (e.g. bB concerning reachability in (16)), we use an uppercase letter for the current block $B(i)$ and a lowercase for the preceding block B(i1)B(i-1). For example, the factor set B4 is the set of positions corresponding to the factor $0^{m-\alpha}$ in $B(i)$, while b4 is the set of positions corresponding to the factor $0^{m-\alpha^{\prime}}$ in B(i1)B(i-1); throughout the tables, we write $\alpha^{\prime}=\alpha(i-1)$, $\beta^{\prime}=\beta(i-1)$ and $\gamma^{\prime}=\gamma(i-1)$. In the reachability tables, distances are measured from the beginning of the relevant factor to the end of the preceding block. For example, the distance to the end of the $B$-block of the factor b4 is $d(\texttt{b4})=m-\gamma^{\prime}+\alpha^{\prime}=m-k$ (by Lemma 7), and its length is $|\texttt{b4}|=\gamma(i-1)=\gamma^{\prime}$. Similarly, $d(\texttt{c4})=2m+k\text{ and }d(\texttt{c8})=m+\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}\text{ and }|\texttt{c8}|=m.$
The value of a factor set is its factor-value. A factor-option of $F$ is a factor $F^{\prime}$ such that there is an $x\in\texttt{F'}$ such that, for some $y\in\texttt{F}$, $x-y\in\{m,2m+k,3m+k\}=\{s_{1},s_{2},s_{3}\}$. To read the tables: fix a column factor set. Each row lists all potential factor-options of a relevant game value. Reachability tables ensure the existence of smaller values as move options, while anti-collision tables exclude equal values in the set of move options; here a triple of inequalities corresponds to the moves $s_{1},s_{2},s_{3}$, respectively.
Example A3 in sinkB vs. bZB. This is the case where $\beta(i-1)=\alpha(i)=0$, and hence $\zeta(i-1)=0^{m}$. Moreover, it follows that $\alpha(i-1)=m-k$. The factor set A3 has size $m$ in such $B(i)$, and, by (16), we study the word
@@ 3 Final remarks
On a more philosophical matter, with respect to the simplicity of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games. However this type of result has been on the list of conjectures for a long time, and our conjectured reciprocity with the wall convention requires further research. For example, Richard Guy writes in the first problem, in the first edition of Unsolved problems in combinatorial games [5] “Subtraction games are known to be periodic. Investigate the relationship between the subtraction set and the length and structure of the period.”; see also [7]. The second part of the conjecture for three move games [3, 8] concerns the non-additive three-move rulesets. Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [3], and has not yet been fully understood. See also the recent survey on subtraction games [6].
Acknowledgements. We thank Richard Nowakowski for pointing out the reference [1]. The images in Figure 1 were generated with assistance from ChatGPT 5.2.